OpenStudy (anonymous):

20) Use Newton's method to find the negative fourth root of 5 by solving the equation x^4 - 5 = 0. Start with x1 = -1 and find x2.

OpenStudy (tkhunny):

\(f(x) = x^{4} - 5\) \(f'(x) = 4x^{3}\) \(x_{n+1} = x_{n} - \dfrac{f(x_{n})}{f'(x_{n})}\) Go!!

OpenStudy (tkhunny):

You can do it. You have \(x_{1} = -1\). I defined \(f(x)\;and\;f'(x)\). Calculate \(f(x_{1})\;and\;f'(x_{1})\) Then what?

OpenStudy (anonymous):

f'(-1)=-4

OpenStudy (tkhunny):

Please don't change your mind after you are already correct. No guessing. \(f(-1) = (-1)^{4} - 5 = 1 - 5 = -4\)

OpenStudy (tkhunny):

Now, build \(x_{2}\).

OpenStudy (anonymous):

how do i

OpenStudy (tkhunny):

Seriously? If you can't read a formula, you are going to suffer in this course. Go look at my first post, again, and use n = 1.

OpenStudy (tkhunny):

\(x_{1} = −1 \) \(f(−1) = −4 \) \(f′(−1) = −4 \) \(x 2 =−1−\dfrac{-4}{-4} = \) ?? Use the formula. Don't make it up.

OpenStudy (anonymous):

x2=-2

OpenStudy (anonymous):

so do i need to do something with -2?

OpenStudy (tkhunny):

Awesome. Now find \(f(x_{2})\;and\;f'(x_{2})\) and do it all again to find \(x_{3}\).

OpenStudy (anonymous):

but is that part of solving the problem?

OpenStudy (tkhunny):

??? What does this mean to you? \(x_{n+1} = g(x_{n})\)? It is an iterative process. You just keep doing it! That's why one index leads to the next.

OpenStudy (anonymous):

f(-2)=-21 f'(-2)=-32 im clueless to how many iteration are needed

OpenStudy (tkhunny):

Did you miss the part where I said, "You just keep doing it"? It is an iterative process. Each iteration should bring you closer to the solution. You quit when you believe you are close enough. Now go! Iterate!

OpenStudy (anonymous):

but i found x2 right? -2

OpenStudy (tkhunny):

Now go find \(x_{3}\).

OpenStudy (anonymous):

im trying to figure what to sub into equation

OpenStudy (tkhunny):

Why? Substitute \(x_{2}\). What other candidate is there?

OpenStudy (anonymous):

-2-1=-2-(-21/-32)

OpenStudy (tkhunny):

What is that "-2-1" on the left-hand side? That should say \(x_{3}\).

OpenStudy (anonymous):

i followed the formula

OpenStudy (tkhunny):

No. If you had followed the formula, with n = 2, the left hand side is \(x_{3}\). Also, \(f(-2) = 11 \ne -21\)

OpenStudy (anonymous):

-16-5=-21

OpenStudy (anonymous):

where is this n coming from?

OpenStudy (anonymous):

-2(-2)=4 4(-2)=-8 -8(-2)=16

OpenStudy (tkhunny):

Is was n = 1 when we started. This leads to 1+1 = 2. This leads to 2+1 = 3. This leads to 3+1 = 4. etc. This is why \(n_{1} = -1\) was supplied in the beginning.

OpenStudy (anonymous):

u started using n but it was x1

OpenStudy (anonymous):

-53/32

OpenStudy (tkhunny):

You don't seem to know what an index is. \(x_{1} = -1\) \(x_{2} = x_{1} - \dfrac{f(x_{1})}{f'(x_{1})} = -1 - \dfrac{-4}{-4} = -1 - 1 = -2\) \(x_{3} = x_{2} - \dfrac{f(x_{2})}{f'(x_{2})} = -2 - \dfrac{11}{-32} = -1.65625\) \(x_{4} = x_{3} - \dfrac{f(x_{3})}{f'(x_{3})} = -1.65625 - \dfrac{2.5249490737915}{-18.1734619140625} = -1.51731394666402\) Are we seeing it, yet? The index is just a counter. There is no evaluation of the index. It is just counting for us.

OpenStudy (tkhunny):

Yes! \(x_{3} = -53/32 = -1.65625\)

OpenStudy (tkhunny):

Now find \(x_{4}\)

OpenStudy (anonymous):

well it seems you have given x4 even though all the numbers are confusing

OpenStudy (anonymous):

so i have to find the negative fourth root but i have to fin x2